t requires two sticks of butter to make a batch of 24 cookies. How much butter will it take to make 150 



1
Expert's answer
2021-02-25T03:44:22-0500

If t requires two sticks of butter to make 24 cookies, one can see that you need 2/24 = 1/12 (one twelve) of the stick of butter to make 1 cookie. Therefore, in order to make 150 cookies, you will need "\frac{1}{12}" ×150 = 12.5 sticks of butter.